OpenStudy (anonymous):

is the solution to 9 to the square root -3 or 3. or maybe both?

OpenStudy (anonymous):

what do you mean exactly?

OpenStudy (mathstudent55):

Do you mean "what is the square root of 9?"

OpenStudy (mathstudent55):

Is it 3 or -3?

OpenStudy (anonymous):

yes i mean what is the square root of 9

OpenStudy (anonymous):

sqrt(9) = 3 -sqrt(9) = -3

OpenStudy (anonymous):

and if it is 3 or -3

OpenStudy (anonymous):

but -3 times -3 is 9

OpenStudy (anonymous):

so the square root of 9 can be -3, right?

OpenStudy (anonymous):

yes, but √ means positive root, not negative.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

if you want it to be negative, you must put the negative sign in the front. I.e. -√(9) = -3

OpenStudy (anonymous):

but where i learned it, the teacher said that to find the square root you had to solve the equation, a times a equals 9

OpenStudy (anonymous):

and -3 times -3 equals 9

OpenStudy (anonymous):

that's a different question than just asking what is √(9) if you have an equation, like x^2 = 9 then when you take the square for both sides, you're actually taking both positive and negative root. I.e x = +/- √(9) which means x = √(9) or x = -√(9) x = 3 or x = -3 now, that does not mean √(9) = +/- 3. If you read carefully, it means √(9) = 3 -√(9) = -3

OpenStudy (anonymous):

i get it now Thanks!